# NormalFlow: Fast, Robust, and Accurate Contact-based Object 6DoF Pose Tracking with Vision-based Tactile Sensors

### v1

@@ Abstract
inline,color=red!40]Wenzhen: I don’t like pushing people to check the video in the paper, especially in the abstract. A paper should be self-explanatory without the video.
@@ III Method
inline,color=red!40]Wenzhen: You should add some demonstration figures for this section We aim to use vision-based tactile sensors to track the 6DoF object poses relative to the sensor’s coordinate frame during contact, without requiring the objects’ geometrical models. Utilizing photometric stereo methods, surface normal maps are derived from the stream of tactile images [16]. The key insight of our NormalFlow algorithm is to find the transformation by directly minimizing discrepancies between normal maps rather than point clouds. This method prevents the accumulated noise associated with the integration process when converting normal maps to point clouds. In contrast, existing approaches [5] [12] used ICP or FilterReg [18] to register point clouds and are prone to noise accumulation.
@@ IV-A5 Long-horizon Tracking Results
Figure 8: Long-horizon tracking results. For each column: [top] sensor trajectories with poses as RGB coordinate axes; [bottom] true 6DoF sensor movements tracked by MoCap (solid lines) and NormalFlow estimation error (shaded area), which is often too small to be seen.inline,color=red!40]Wenzhen: The frames are hard to read. YOu might want to get the drawing sparser. inline,color=red!40]Wenzhen: I’m confused about the way you draw the result plots: why do you use shades to indicate error, instead of just drawing the ”ground truth” plot and the ”measured” plot?inline,color=blue!20]Joe: I tried that, and it looks very messy since they are too close. I also tried other visualization approaches, such as error bars, draw only markers, and I think this version is the best.

### v2

@@ III Method
We aim to use vision-based tactile sensors to track the 6DoF object poses relative to the sensor’s coordinate frame during contact, without requiring the objects’ geometrical models. Utilizing photometric stereo methods, surface normal maps are derived from the stream of tactile images [16]. The key insight of our NormalFlow algorithm is to find the transformation by directly minimizing discrepancies between normal maps rather than point clouds. This method prevents the accumulated noise associated with the integration process when converting normal maps to point clouds. In contrast, existing approaches [5] [12] used ICP or FilterReg [18] to register point clouds and are prone to noise accumulation.
@@ IV-A5 Long-horizon Tracking Results
Figure 8: Long-horizon tracking results. For each column: [top] sensor trajectories with poses as RGB coordinate axes; [bottom] true 6DoF sensor movements tracked by MoCap (solid lines) and NormalFlow estimation error (shaded area), which is often too small to be seen.